OpenStudy (anonymous):

Write the equation of the line in slop-intercept form that passes through the points (-3,2) and (5,3).

OpenStudy (austinl):

y=mx+b \[m = \frac{y_2-y_1}{x_2-x_1}\] Then plug in a point and solve for b.

OpenStudy (anonymous):

ok thanks very much

OpenStudy (austinl):

No problem!

OpenStudy (anonymous):

Graph the line 2x+3y=6

OpenStudy (anonymous):

Some please help me!!! Graph the line 2x+3y=6